OpenStudy (anonymous):

find the x- and y- intercepts 2x - 3y = 9

OpenStudy (anonymous):

For the \(x\)-intercepts, we simply plug in 0 for \(y\) and, similarly do the opposite for the \(y\)-intercept.

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

so if you want to solve for the x intercept of the equation you would substitute y for 0 so that it becomes 2x-3(0)=9

OpenStudy (anonymous):

typo mb, anyway if you want to solve for the y-intercept you would substitute x for 0 so that it becomes 2(0)-3y=9 -3y=9 y=-3

OpenStudy (anonymous):

then the next one should be 2x - 3(0) = 9 2x=9 x= 9/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

np